# How Far Are We from Believable AI Agents? A Framework for Evaluating the Believability of Human Behavior Simulation

### v1

@@ 7 Conclusion
We propose two novel metrics to measure agents’ believability level: long consistency and robustness. We introduce SimulateBench: a benchmark for the agents to simulate characters and measure agents’ consistency and robustness. Through the SimulateBench, we evaluate the level of believability of agents composed of popular LLMs. We hope that our work inspires the evaluation of simulated agents in the future.
@@ A.1 Profile Descriptive Framework
  • Social Role. Social role (Wasserman and Faust, 1994; Eagly and Wood, 2012) refers to a set of connected behaviors, obligations, beliefs, and norms as conceptualized by people in a social situation. We will record the characters’ roles in different social situations. Besides, drawing inspiration from Dunbar et al. (1997); Gao et al. (2023b), we also document the role’s traits, routines/habits, general experiences, and plans/goals to enhance LLMs’ simulation performance in social interactions.
  • Relationship. In the context of social interactions, the relationship can influence the agent’s response in a discussion, the actions to be taken, the willingness to collaborate, and their inclination to diffuse information. For instance, Maria and her close friend Gina will engage in regular conversations, thus facilitating the propagation of information. Hence, in order to facilitate the agent’s simulation of behaviors that align with the relationship between the agent and others in social interaction, we will comprehensively document the following aspects of the relationship: familiarity, judgment, affection, behavioral patterns, relationship status, and communication history.
@@ A.3 Character Variants Dataset
  • Surname Inspired by Aher et al. (2023b), we will replace the surname of the character Homer in The The Simpsons to investigate whether the LLMs’ simulated performance will be influenced. Aher et al. (2023b) have listed the most common surnames in each of the five races. Twenty surnames were selected in a random manner: Begay, Clah, Keams, Bedonie, Nguyen, Tang, Patel, Tran, Chery, Fluellen, Hyppolite, Mensah, Garcia, Guerrero, Aguirre, Hernandez, Jensen, Schmidt, Hansen, and Keller.
@@
Table 8: The version and context size of LLMs evaluated in our work
@@ C.1 Evaluated Models
We assess the believability of 10 LLMs, their release time and context size are listed in Table 8.

### v2

@@ 7 Conclusion
We proposed two novel dimensions to measure LLMs’ level of believability: consistency and robustness. We introduced SimulateBench, a benchmark for the profile collection and the measure of LLMs’ consistency and robustness. Through the SimulateBench, we evaluated the level of believability of popular LLMs. Our experimental results and findings provided insights to facilitate future research on developing human-like AI.
8 Limitations
In this paper, we proposed two dimensions to measure LLMs’ level of believability when simulating human behavior. Simulating human behavior is an intricate undertaking that necessitates extensive and detailed information on the character’s profile. Despite the fact that our work has a considerably thorough profile compared to earlier works, it may still be inadequate. Furthermore, despite our thorough evaluation of many well-known models, certain commercial models, such as Claude from Anthropic, have not been included in our evaluation. This omission is due to the requirement of qualification audits for using these models, which we do not have access to. Consequently, the evaluation of these models is not included in our research.
9 Ethics Statement
We strictly adhere to the ACL Code of Ethics. We placed high importance on ensuring the comfort and well-being of our annotators. We advised them to stop the annotation process if they came across any information that caused them discomfort. We utilize the models in accordance with their designated purpose. We recruited annotators at a rate of 2 $\sim$ 3 times their local hourly minimum wage. According to the annotators’ feedback, they spent approximately 30 hours collecting profile data and questions for a character. In summary, we make every effort to adhere to the ethical norms set forth by ACL.
@@ A.1 Profile Descriptive Framework
  • Social Role. Social role (Wasserman, 1994; Eagly and Wood, 2012) refers to a set of connected behaviors, obligations, beliefs, and norms as conceptualized by people in a social situation. We will record the characters’ roles in different social situations. Furthermore, drawing inspiration from Dunbar et al. (1997); Gao et al. (2023), we also document the role’s traits, routines/habits, general experiences, and plans/goals to enhance LLMs’ simulation performance in social interactions.
  • Relationship. In the context of social interactions, the relationship can influence the LLMs’ response in a discussion, the actions to be taken, the willingness to collaborate, and their inclination to diffuse information. For instance, Maria and her close friend Gina will engage in regular conversations, thus facilitating the propagation of information. Hence, in order to facilitate the LLMs’ simulation of behaviors that align with the relationship between the LLM and others in social interaction, we will comprehensively document the following aspects of the relationship: familiarity, judgment, affection, behavioral patterns, relationship status, and communication history.
@@ A.3 Profile perturbations
  • Surname Inspired by Aher et al. (2023), we will replace the surname of the character Homer in The The Simpsons to investigate whether the LLMs’ simulated performance will be influenced. Aher et al. (2023) have listed the most common surnames in each of the five races. Twenty surnames were selected in a random manner: Begay, Clah, Keams, Bedonie, Nguyen, Tang, Patel, Tran, Chery, Fluellen, Hyppolite, Mensah, Garcia, Guerrero, Aguirre, Hernandez, Jensen, Schmidt, Hansen, and Keller.
@@
A.4 Question generation
Prompt used to generate question about immutable characteristic
"I need your expertise in questionnaire design. I want you to create a set of one-choice questions that will gather basic information about a person. Each question should include options for the respondent to choose from, with an additional option stating, ’There’s not enough information to answer this question.’ Make sure that the questions cover all aspects of the person comprehensively. Remember, the goal is to obtain detailed and accurate responses. Please avoid imposing any assumptions or biases in your questions."
Prompt used to generate question about social role
"I need your expertise in questionnaire design. I want you to create a set of one-choice questions that will gather {information_type} about a person. Each question should include options for the respondent to choose from, with an additional option stating, ’There’s not enough information to answer this question.’ Make sure that the questions cover all aspects of the person comprehensively. Remember, the goal is to obtain detailed and accurate responses. Please avoid imposing any assumptions or biases in your questions."
Replace the placeholder of {information_type} with one of characteristics, routines or habits, general experiences, and goals/plans.
Prompt used to generate question about relationship
"I need your expertise in questionnaire design. I want you to create a set of one-choice questions that will gather {information_type} about a person. Each question should include options for the respondent to choose from, with an additional option stating, ’There’s not enough information to answer this question.’ Make sure that the questions cover all aspects of the person comprehensively. Remember, the goal is to obtain detailed and accurate responses. Please avoid imposing any assumptions or biases in your questions."
Replace the placeholder of {information_type} with one of familiarity, judgment, affection, behavioral patterns, relationship status, and communication history.
Example of questions
The questions in the consistency dataset are categorized into two classes according to whether there is insufficient information about the character in the profile for the LLM to deduce the answer. Examples are listed in Figure 3.
Figure 3: An illustrative example of the question types of Known and Unknown.
@@ C.1 Evaluated Models
We assess the believability of 10 LLMs; their release time and context size are listed in Table 7.
Table 7: The version and context size of LLMs evaluated in our work
Appendix D Details for Influential Factors of Believability
D.1 Examine the Effect of Simulation hallucination
A possible explanation of simulation hallucination is that the model might have known the answer to a question due to the knowledge learned in the training process, even if the answer is not in the agent’s profile, so the model prefers to answer the question rather than answer with "I do not know." as required in the prompt. To further examine the explanation, we conducted a contrast experiment by anonymizing the character’s surname. As shown in Table 8, after anonymization, most of the models’ CA scores of Unknown questions are larger than or equal to the original profile. Some cases where the GPT-3.5-Turbo-16K correctly answers the Unknown question after anonymization are shown in Figure 4.
Models Immutable Characteristic Original Keams Bedonie Nguyen GPT-3.5-Turbo-16K 0.58 0.74 0.79 0.63 ChatGLM2-6B-32K 0.53 0.58 0.58 0.63 XVERSE-13B-Chat 0.47 0.47 0.47 0.53 Vicuna-7B-16K 0.32 0.37 0.37 0.37 GPT-4 0.21 0.16 0.16 0.16 ChatGLM2-6B 0.21 0.21 0.21 0.21 Vicuna-13B-16K 0.11 0.00 0.16 0.00 Qwen-14B-Chat 0.11 0.11 0.11 0.11 Qwen-7B-Chat 0.00 0.00 0.00 0.05 LongChat-7B-32K 0.00 0.00 0.00 0.00
Table 8: The CA scores of ten models to answer the Unknown questions of Immutable Characteristic. The Original refers to the character’s profile being unchanged. Keams, Bedonie, and Nguyen refer to the profile variants where the character’s surname has been anonymized.
Figure 4: Cases where GPT-3.5-Turbo-16K answer the Unknown questions correctly after anonymization.